log x/2a+3b-5c = log y/2b+3c-5a = log z/2c+3a-5b , then xyz=?? Show the steps..

Grade:

1 Answers

Arun
5 years ago
Dear student
 
logx/2a+3b-5c=logy/2b+3c-5a=logx/2c+3a-5bTherefore, we can write it as logx+logy+ logz/0So, logx+ logy+ logz=0So,log(xyz)=0 , xyz=1
 
Regards
Arun